OpenStudy (anonymous):

Solve the following problems using the sequence: five, thirteen, twenty-one, twenty-nine, thirty-seven, and so on a. Use words to write a rule for the sequence. b. Use the rule you wrote in the previous part to find the next three terms of the sequence. Be sure to show all your steps.

OpenStudy (anonymous):

write the numbers as numbers rather than words, a pattern will emerge

OpenStudy (kewlgeek555):

Well think about it. First of all, you need to know how to write a rule for a function. Now, I forgot how to write rules but I think it was something like "Add x" or something like that. Now think about it. from 5 to 13 is 8 and from 13 to 21 is 8 and from 21 to 29 is 8! So the rule would be "Add 8" Hope this helps for the future.

OpenStudy (anonymous):

y=x+8 Next=now+8